Use the table below to answer the questions that follow. Define productivity as GDP per worker, and please enter all answers as numerical values rounded to 3 decimal places, and not as percentages (i.e. 0.103 instead of 10.3%).

GDP Population Workers
Year = 1 1027 155 129
Year = 2 1247 279 233

1) What is the GDP growth rate?

2) What is GDP per capita growth rate?

3) What is the productivity growth rate?
Answer #1

GDP growth rate = Increase in value / Original value = (1247 - 1027)/1027 = 0.214

GDP per capita = GDP / Population. GDP per capita growth rate = (1247/279 - 1027/155)/(1027/155) = -0.325

Productivity = GDP/ workers. Productivity growth rate = (1247/233 - 1027/129)/(1027/129) = -0.328